A researcher records the amount of time (in minutes) that parent-child pairs spent on social networking sites to test whether they show any generational differences. From the following findings reported in APA format, interpret these results.

Parents spent significantly less time on social networking sites compared to their children

(MD = 42 minutes),

t(19) = 3.828, p < 0.05, d = 0.51.


State the research design used (repeated measures or matched pairs).

repeated measures

OR

design matched pairs design    


State the sample size.

______ participants

State the decision to retain or reject the null hypothesis.

Retain the null hypothesis.

OR

Reject the null hypothesis.    


State the effect size.

small

OR

medium

OR

large
Answer #1

i) The research design used : design matched pairs design   

ii) The test statistic can be written as

n * Sd which under H0 follows a t distribution with n-1 df.

We reject H0 at 5% level of significance if tobst0.05,n-1 or p-value < 0.05

critical value = t0.05,n-1

sample size = n = (n -1 ) +1 = 19 + 1 = 20

Since p-value < 0.05, so we reject the null hypothesis.

ans-> effect size : large